OpenStudy (anonymous):

Metal award ^^ A man covers d miles in t hrs. At that rate how long (in hrs) will in take him to cover m miles?

OpenStudy (blurbendy):

depends if m miles is > or < d miles

OpenStudy (anonymous):

okidoki!! ^^ Thanks.

OpenStudy (blurbendy):

np!

OpenStudy (blurbendy):

if you wanted to be more specific you could say r = d/t

OpenStudy (blurbendy):

since d = rt

OpenStudy (blurbendy):

so t = d/r

OpenStudy (anonymous):

since distance = rate x time, then time = distance/rate so the time you want is: time = \(\ \frac{m}{d/t} \)